22. One side of a parallelogram exceeds the other by 30cm. If the perimeter of the parallelogram is 260cm, find the lengths of the sides of the parallelogram.
Answer by psbhowmick(878)

Let the two sides of the parallelogram be x and y.

2(x+y)=260
x+y=130 ______ (1)

Also, x-y=30 ______ (2)

Adding (1) and (2)
2x = 160
x = 80
y = 130 - 80 = 50

So the sides of the parallelogram are 80 cm and 50 cm.